OpenStudy (anonymous):

Find a polynomial of least possible degree. Use the leading coefficient of 1 or -1.

OpenStudy (anonymous):

to me, the least possible degree of a polynomial is 1 .and that can have a form of ax +b and if you choose coefficient 1 or -1 , just plug them into the form P(x) = x +1 or P(x)= -x -1 or P(x)= x -1. Hopefully mine is not wrong